OpenStudy (therealmeeeee):

plzzzz help medal given Simplify the following expression. 4 - (8 × 4)10 - 26 -342 298 -290 516

OpenStudy (anonymous):

-342 lol

OpenStudy (therealmeeeee):

Simplify the following expression. 3 × 3 + 4 ÷ 2 - (5 + 1) 7 5 13 17

OpenStudy (solomonzelman):

\[4 - (8 × 4)10 - 26\]\[4 - (32)10 - 26\]\[4 - 320 - 26=?\]right 342

OpenStudy (solomonzelman):

i meant -342.

OpenStudy (therealmeeeee):

Simplify the following expression. 3 × 3 + 4 ÷ 2 - (5 + 1) 7 5 13 17

OpenStudy (therealmeeeee):

i got 5

OpenStudy (therealmeeeee):

this wat i did Order of Operations: 1. Parentheses or brackets from the inside out 2. Multiplication and division from left to right 3. Addition and subtraction from left to right 3 × 3 + 4 ÷ 2 - (5 + 1) = 3 × 3 + 4 ÷ 2 - 6 = 9 + 4 ÷ 2 - 6 = 9 + 2 - 6 = 11 - 6 = 5

OpenStudy (anonymous):

I agree with you, @TheRealMeeeee